OpenStudy (anonymous):

What is an easy process to solve this please! y <= x - 5 y <= -x - 4

OpenStudy (anonymous):

I will check the link out!

OpenStudy (anonymous):

do you still need help with this @proverbsm

OpenStudy (anonymous):

yerp ;l

OpenStudy (jdoe0001):

so you have y =x - 5 y =-x - 4 I assume so, just grab the top "y" and plug it in the second, as the link example showed what does it give you?

OpenStudy (anonymous):

is it y=1? lmao I'm so lost. This sucks!

OpenStudy (jdoe0001):

well, what's "y" equals to in the 1st equation?

OpenStudy (anonymous):

x-5

OpenStudy (jdoe0001):

so grab THAT and stick it in the 2nd equation instead of "y" how would the 2nd equation would look like?

OpenStudy (anonymous):

x-5-x-4 , uhm maybe? <--

OpenStudy (jdoe0001):

$$ y =\color{red}{x - 5}\\ y =-x - 4\\ \text{thus}\\ \color{red}{x-5}=-x - 4 $$

OpenStudy (anonymous):

i had that but wasnt ve

OpenStudy (anonymous):

very sure of it

OpenStudy (anonymous):

okay now that I know what to do with my equation, I solve it from here?

OpenStudy (jdoe0001):

see once you have THAT only, then notice, the "y" is really gone, and all you'ref left with is "x's" so, now just solve for "x"

OpenStudy (anonymous):